# FakeShield: Explainable Image Forgery Detection and Localization via Multi-modal Large Language Models

### v3

@@ 3.5 Training Objectives
The two submodules of our FakeShield are trained end-to-end separately. For DTE-FDM, the domain tag generator utilizes cross-entropy loss $\ell_{ce}$ as its training objective, enabling it to distinguish between different data domains. Following the approach of LLaVA (Liu et al., 2023), our LLM’s training objective is the autoregressive text generation cross-entropy loss. The training target of DTE-FDM $\ell_{det}$ can be formulated as:
$\ell_{det}=\ell_{ce}(\hat{\mathbf{O}}_{det},\mathbf{O}_{txt})+\lambda\cdot\ell_{ce}(\hat{\mathbf{T}}_{tag},\mathbf{T}_{tag}),$ (4)
where $\ell_{ce}$ represents cross-entropy loss, and $\lambda$ denotes the weight used for balancing different loss components. For MFLM, we apply $\ell_{ce}$ to constrain TCM to produce high-quality prompt $\hat{\mathbf{y}}_{txt}$ with <SEG> token. Meanwhile, we use a linear combination of binary cross-entropy loss $\ell_{bce}$ and dice loss $\ell_{dice}$ to encourage the output of MFLM $\hat{\mathbf{M}}_{loc}$ to be close to the GT mask $\mathbf{M}_{loc}$. Given the ground-truth prompt $\mathbf{y}_{txt}$ (such as “It is <seg>”) and mask $\mathbf{M}_{loc}$, our training losses for MFLM $\ell_{loc}$ can be formulated as:
@@ 4.1 Experimental Setup
Dataset: We employ the dataset construction method outlined in Section 3.1 to build the training and test sets of the MMTD-Set. For the training set, we utilize PhotoShop tampering (e.g., CASIAv2 (Dong et al., 2013), Fantastic Reality (Kniaz et al., 2019)), DeepFake tampering (e.g., FFHQ, FaceApp (Dang et al., 2020)), and some self-constructed AIGC-Editing tampered data as the source dataset. For the testing set, we select several challenging public benchmark datasets including PhotoShop tampering (CASIAv1+ (Dong et al., 2013), Columbia (Ng et al., 2009), IMD2020 (Novozamsky et al., 2020), Coverage (Wen et al., 2016), DSO (De Carvalho et al., 2013), Korus (Korus & Huang, 2016)), DeepFake tampering (e.g., FFHQ, FaceApp (Dang et al., 2020)), and some self-generated AIGC-Editing data. See the Appendix for more details.
State-of-the-Art Methods: To ensure a fair comparison, we select competitive methods that provide either open-source code or pre-trained models. To evaluate the IFDL performance of FakeShield, we compare it against SPAN (Hu et al., 2020), MantraNet (Wu et al., 2019), OSN (Wu et al., 2022), HiFi-Net (Guo et al., 2023), PSCC-Net (Liu et al., 2022), CAT-Net (Kwon et al., 2021), and MVSS-Net (Dong et al., 2022), all of which are retrained on the MMTD-Set for consistency with the same training setup. For DeepFake detection, CADDM (Dong et al., 2023) and HiFi-DeepFake (Guo et al., 2023) are chosen as comparison methods. Additionally, to assess the explanation ability of FakeShield, we compare it with open-source M-LLMs such as LLaVA-v1.6-34B (Liu et al., 2023), InternVL2-26B (Chen et al., 2024), and Qwen2-VL-7B (Wang et al., 2024), as well as the closed-source model GPT-4o (OpenAI, 2023).
Evaluation Metrics: For detection, we report image-level accuracy (ACC) and F1 scores. We use Intersection over Union (IoU) and F1 scores for localization. To evaluate interpretability, we use Cosine Semantic Similarity (CSS) to assess the similarity between the predicted text and ground truth text by calculating the cosine similarity between their high-dimensional semantic vectors. For the IFDL, a default threshold of $0.5$ is applied unless otherwise specified.
Implementation Details: On the MMTD-Set, we initially fine-tune the M-LLM using LoRA (rank=$128$, alpha=$256$), such as LLaVA-v1.5-13B (Liu et al., 2023), while simultaneously training the Domain Tag Generator with full parameters. The model is trained for $10$ epochs on $4$ NVIDIA A100 40G GPUs, with a learning rate of $2$$\times$$10^{-4}$ and a batch size of $24$. Afterward, we fine-tune the Tamper Comprehension Module and Segment Anything Model (Kirillov et al., 2023) with LoRA (rank=$8$, alpha=$16$), training for $24$ epochs on the same hardware configuration, with a learning rate of $3$$\times$$10^{-4}$ and a batch size of $48$.
@@ 4.2 Comparison with Image Forgery Detection Method
Table 2: Performance comparison with competitive DeepFake detection methods.
Method CADDM HiFi-DeepFake Ours
ACC 0.5227 0.5177 0.9835
F1 0.5982 0.6403 0.9915
Furthermore, we compare our method with several state-of-the-art DeepFake detection approaches on the MMTD-Set(DeepFake), including CADDM and HiFi-DeepFake. As shown in Table 2, our approach significantly outperforms the other two methods. Notably, the domain tag mechanism we introduce not only alleviates conflicts across diverse data domains but also promotes complementarity and mutual enhancement between them. Despite being specifically designed for DeepFake detection, these comparison methods still fall short of the performance achieved by FakeShield.
@@ 4.3 Comparison with M-LLMs
To assess the quality of explanation text generation, we employ the tampered area descriptions generated by GPT-4o as ground truth on the MMTD-Set(Photoshop, DeepFake, AIGC-Editing), using cosine semantic similarity (CSS) to compare the performance of pre-trained M-LLMs against FakeShield. The results, presented in Table 3, demonstrate that our approach consistently achieves the best performance across nearly all test sets. For instance, on the DSO, our method attains a CSS score of 0.8873, significantly surpassing the second-best result from InternVL2-26B, which only achieves 0.6484. A selection of model outputs is illustrated in the Appendix.
It is important to highlight that some pre-trained M-LLMs still exhibit a degree of proficiency in detecting tampered content. For instance, in cases involving blatant violations of physical laws due to tampering, these M-LLMs leverage their pre-training knowledge to make reasonably correct judgments. However, they struggle to perform more precise analyses, such as identifying inconsistencies in lighting or perspective, leading to lower overall accuracy.
@@ 4.4 Comparison with Image Forgery Location Method
To assess the model’s capability to locate tampered regions, we conduct comparisons to establish some competitive IFDL methods on MMTD-Set (Photoshop, AIGC-Editing). As present in Table 4, our method consistently surpasses the performance of others across almost all test datasets. For instance, on the IMD2020, our method outperforms the suboptimal method, OSN, with notable advantages of 0.12 in IoU and 0.1 in F1 score. On the CASIA1+, we also lead the sub-optimal method OSN with an IoU of 0.07 and an F1 of 0.09.
Additionally, subjective comparisons of several methods are illustrated in Figure 4, where it is evident that our approach precisely captures the boundaries of tampered areas, producing clean and complete segmentations. In contrast, methods like PSCC-Net exhibit dispersed attention over the image, resulting in blurred segmentations and an overly broad predicted tampering range. Notably, as our segmentation module MFLM is based on the pre-trained visual segmentation model SAM, it inherits SAM’s powerful semantic segmentation capabilities and can accurately segment targets with clear semantic information (Columns 1 and 2 of Figure 4). Additionally, due to the diverse tampering types in MMTD-Set, our method is also capable of accurately segmenting tampered areas that lack distinct semantic information (Columns 3 of Figure 4).
Table 4: Comparative results of tamper localization capabilities between competing IFDL methods and FakeShield, tested on MMTD-Set(Photoshop, AIGC-Editing).FakeShield consistently achieves the highest IoU and F1 scores across most test sets.
Method CASIA1+ IMD2020 Columbia NIST DSO Korus AIGC-Editing
IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
SPAN 0.11 0.14 0.09 0.14 0.14 0.20 0.16 0.21 0.14 0.24 0.06 0.10 0.09 0.12
ManTraNet 0.09 0.13 0.10 0.16 0.04 0.07 0.14 0.20 0.08 0.13 0.02 0.05 0.07 0.12
OSN 0.47 0.51 0.38 0.47 0.58 0.69 0.25 0.33 0.32 0.45 0.14 0.19 0.07 0.09
HiFi-Net 0.13 0.18 0.09 0.14 0.06 0.11 0.09 0.13 0.18 0.29 0.01 0.02 0.13 0.22
PSCC-Net 0.36 0.46 0.22 0.32 0.64 0.74 0.18 0.26 0.22 0.33 0.15 0.22 0.10 0.15
CAT-Net 0.44 0.51 0.14 0.19 0.08 0.13 0.14 0.19 0.06 0.10 0.04 0.06 0.03 0.05
MVSS-Net 0.40 0.48 0.23 0.31 0.48 0.61 0.24 0.29 0.23 0.34 0.12 0.17 0.18 0.24
FakeShield 0.54 0.60 0.50 0.57 0.67 0.75 0.32 0.37 0.48 0.52 0.17 0.20 0.18 0.24
Figure 4: Testing results on MMTD-Set for FakeShield are compared with PSCC-Net, OSN, CAT-Net, and MVSS-Net. The test samples, from left to right, are drawn from IMD2020, CASIA1+, Columbia, NIST16, Korus, and DSO.
@@ 4.5 Robustness Study
Table 5: Explanation and location performance of our FakeShield under different degradations.
@@
Figure 5: Ablation studies on the LLM in the DTE-FDM.
@@ 4.6 Ablation Study
Ablation Study on Domain Tag Generator: To validate that the proposed domain tag effectively mitigates domain discrepancies and enhances the model’s generalization across diverse data domains, we conducted an ablation study. Specifically, we removed the domain tag generator (DTG) and trained the FakeShield with identical configurations on the MMTD-Set, the test results are displayed in Table 6. Without the DTG module, the model’s detection performance declined across test sets from each data domain. Notably, the detection ACC and F1 score for DeepFake decreased by 0.09. This demonstrates that without the support of the DTG module, the model struggles to effectively differentiate between various data domains, leading to more pronounced data conflicts and a significant reduction in both generalization and practical applicability.
@@
Method CASIA1+ DeepFake AIGC-Editing
ACC F1 ACC F1 ACC F1
Ours w/o DTG 0.92 0.92 0.89 0.90 0.72 0.78
Ours 0.95 0.95 0.98 0.99 0.93 0.93
@@ A.2 Data Sources
A.3 Answer Analysis

### v4

@@ 3.5 Training Objectives
The two submodules of our FakeShield are trained end-to-end separately. For DTE-FDM, the domain tag generator utilizes cross-entropy loss $\ell_{ce}$ as its training objective, enabling it to distinguish between different data domains. Following the approach of LLaVA, our LLM’s training objective is the cross-entropy loss $\ell_{ce}$. The training target of DTE-FDM $\ell_{det}$ can be formulated as:
$\ell_{det}=\ell_{ce}(\hat{\mathbf{O}}_{det},\mathbf{O}_{det})+\lambda\cdot\ell_{ce}(\hat{\mathbf{T}}_{tag},\mathbf{T}_{tag}),$ (4)
where $\lambda$ denotes the weight balancing different loss components, $\mathbf{O}_{det}$ and $\mathbf{T}_{tag}$ represent the predictions of LLM and DTG, while $\hat{\mathbf{O}}_{det}$ and $\hat{\mathbf{T}}_{tag}$ represent their corresponding ground truth. For MFLM, we apply $\ell_{ce}$ to constrain TCM to produce high-quality prompt $\mathbf{y}_{txt}$ with <SEG> token. Meanwhile, we use a linear combination of binary cross-entropy loss $\ell_{bce}$ and dice loss $\ell_{dice}$ to encourage the output of MFLM $\mathbf{M}_{loc}$ to be close to the GT mask $\hat{\mathbf{M}}_{loc}$. Given the ground-truth prompt $\hat{\mathbf{y}}_{txt}$ (e.g., “It is <SEG>”) and mask $\hat{\mathbf{M}}_{loc}$, our training losses for MFLM $\ell_{loc}$ can be formulated as:
@@ 4.1 Experimental Setup
Dataset: We employ the dataset construction method outlined in Section 3.1 to build the training and test sets of the MMTD-Set. For the training set, we utilize PhotoShop tampering (e.g., CASIAv2 (Dong et al., 2013), Fantastic Reality (Kniaz et al., 2019)), DeepFake tampering (e.g., FFHQ, FaceApp (Dang et al., 2020)), and some self-constructed AIGC-Editing tampered data as the source dataset. For the testing set, we select several challenging public benchmark datasets including PhotoShop tampering (CASIA1+ (Dong et al., 2013), Columbia (Ng et al., 2009), IMD2020 (Novozamsky et al., 2020), Coverage (Wen et al., 2016), DSO (De Carvalho et al., 2013), Korus (Korus & Huang, 2016)), DeepFake tampering (e.g., FFHQ, FaceApp (Dang et al., 2020), Seq-DeepFake (Shao et al., 2022)), and some self-generated AIGC-Editing data.
State-of-the-Art Methods: To ensure a fair comparison, we select competitive methods that provide either open-source code or pre-trained models. To evaluate the IFDL performance of FakeShield, we compare it against SPAN (Hu et al., 2020), MantraNet (Wu et al., 2019), OSN (Wu et al., 2022), HiFi-Net (Guo et al., 2023), PSCC-Net (Liu et al., 2022), CAT-Net (Kwon et al., 2021), and MVSS-Net (Dong et al., 2022), all of which are retrained on the MMTD-Set for consistency with the same training setup. For DeepFake detection, CADDM (Dong et al., 2023), HiFi-DeepFake (Guo et al., 2023), RECCE (Cao et al., 2022) and Exposing (Ba et al., 2024) are chosen as comparison methods. Additionally, to assess the explanation ability of FakeShield, we compare it with open-source M-LLMs such as LLaVA-v1.6-34B (Liu et al., 2024), InternVL2-26B (Chen et al., 2024), and Qwen2-VL-7B (Wang et al., 2024), as well as the closed-source model GPT-4o (OpenAI, 2023).
Evaluation Metrics: For detection, we report image-level accuracy (ACC) and F1 scores. For localization, we provide Intersection over Union (IoU) and F1 scores. To evaluate interpretability, we use Cosine Semantic Similarity (CSS) to assess the similarity between the predicted text and ground truth text by calculating the cosine similarity between their high-dimensional semantic vectors. For both detection and localization, a default threshold of $0.5$ is applied unless otherwise specified.
Implementation Details: On the MMTD-Set, we initially fine-tune the M-LLM using LoRA (rank=$128$, alpha=$256$), such as LLaVA-v1.5-13B (Liu et al., 2024), while simultaneously training the Domain Tag Generator with full parameters. The model is trained for $10$ epochs on $4$ NVIDIA A100 40G GPUs, with a learning rate of $2$$\times$$10^{-4}$. Afterward, we fine-tune the Tamper Comprehension Module and Segment Anything Model (Kirillov et al., 2023) with LoRA (rank=$8$, alpha=$16$), training for $24$ epochs on the same hardware configuration, with a learning rate of $3$$\times$$10^{-4}$.
@@ 4.2 Comparison with Image Forgery Detection Method
Table 2: Performance comparison on DFFD and Seq-DeepFake datasets.
Method DFFD Seq-DeepFake
ACC F1 ACC F1
CADDM 0.52 0.60 0.53 0.59
HiFi-DeepFake 0.52 0.64 0.52 0.57
Exposing 0.82 0.84 0.71 0.78
RECCE 0.92 0.92 0.75 0.79
FakeShield 0.98 0.99 0.84 0.91
Furthermore, we compare our method with some recent DeepFake detection approaches on the DFFD (FaceApp, FFHQ) and Seq-DeepFake, where Exposing and RECCE were retrained on our datasets. As reported in Table 2, our approach significantly outperforms all other methods. Notably, the domain tag mechanism not only alleviates conflicts across diverse data domains but also promotes complementarity and mutual enhancement between them. Although specifically designed for DeepFake detection, these methods still underperform compared to our FakeShield.
Figure 4: Detection, localization and explanation results of our FakeShield.
@@ 4.3 Comparison with M-LLMs
To assess the quality of explanation text generation, we employ the tampered descriptions generated by GPT-4o as ground truth on the MMTD-Set(Photoshop, DeepFake, AIGC-Editing), using cosine semantic similarity (CSS) to compare the performance of pre-trained M-LLMs against FakeShield. The results in Table 3 show that our approach consistently achieves the best performance across nearly all test sets. For instance, on the DSO, our method attains a CSS score of 0.8873, significantly surpassing the second-best result from InternVL2-26B. Figure 4 presents an example of our explanation. The model identifies the three airplanes in the sky as fake and explains its judgment based on lighting and edges, providing an accurate manipulation region mask.
Notably, some pre-trained M-LLMs exhibit limited proficiency in detecting tampered content. For instance, when tampering causes clear physical law violations, these M-LLMs leverage their pre-training knowledge to make reasonably correct judgments. However, they struggle to perform more precise analyses, like detecting lighting or perspective inconsistencies, reducing overall accuracy.
@@ 4.4 Comparison with Image Forgery Location Method
To assess the model’s capability to locate tampered regions, we conduct comparisons with some competitive IFDL methods on the MMTD-Set. As present in Table 4, our method consistently surpasses other methods across almost all test datasets. For instance, on the IMD2020, our method outperforms the suboptimal method OSN with notable advantages of 0.12 in IoU and 0.1 in F1 score. On the CASIA1+, we also lead OSN with an IoU of 0.07 and an F1 of 0.09.
Additionally, subjective comparisons of several methods are shown in Figure 5, where our approach precisely captures the tampered areas, producing clean and complete masks. In contrast, methods like PSCC-Net exhibit dispersed attention over the image, resulting in blurred segmentations and an overly broad predicted tampering area. Notably, as our segmentation module MFLM is based on the pre-trained visual model SAM, it inherits SAM’s powerful semantic segmentation capabilities and can accurately segment targets with clear semantic information (Columns 1 and 2 of Figure 5). Additionally, due to the diverse tampering types in MMTD-Set, our method can also accurately segmenting tampered areas that lack distinct semantic information (Columns 3 of Figure 5).
Table 4: Comparative results of tamper localization capabilities between competing IFDL methods and FakeShield, tested on MMTD-Set(Photoshop, DeepFake, AIGC-Editing).
Method CASIA1+ IMD2020 Columbia NIST DSO Korus DeepFake AIGC-Editing
IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1 IoU F1
SPAN 0.11 0.14 0.09 0.14 0.14 0.20 0.16 0.21 0.14 0.24 0.06 0.10 0.04 0.06 0.09 0.12
ManTraNet 0.09 0.13 0.10 0.16 0.04 0.07 0.14 0.20 0.08 0.13 0.02 0.05 0.03 0.05 0.07 0.12
OSN 0.47 0.51 0.38 0.47 0.58 0.69 0.25 0.33 0.32 0.45 0.14 0.19 0.11 0.13 0.07 0.09
HiFi-Net 0.13 0.18 0.09 0.14 0.06 0.11 0.09 0.13 0.18 0.29 0.01 0.02 0.07 0.11 0.13 0.22
PSCC-Net 0.36 0.46 0.22 0.32 0.64 0.74 0.18 0.26 0.22 0.33 0.15 0.22 0.12 0.18 0.10 0.15
CAT-Net 0.44 0.51 0.14 0.19 0.08 0.13 0.14 0.19 0.06 0.10 0.04 0.06 0.10 0.15 0.03 0.05
MVSS-Net 0.40 0.48 0.23 0.31 0.48 0.61 0.24 0.29 0.23 0.34 0.12 0.17 0.10 0.09 0.18 0.24
FakeShield 0.54 0.60 0.50 0.57 0.67 0.75 0.32 0.37 0.48 0.52 0.17 0.20 0.14 0.22 0.18 0.24
Figure 5: Comparisons between our FakeShield and other competitive methods. The samples, from left to right, are drawn from IMD2020, CASIA1+, Columbia, NIST16, Korus, DSO, and DeepFake.
@@ 4.5 Robustness Study
Table 5: Explanation and location performance under different degradations.
@@
Figure 6: Ablation study on the DTE-FDM.
@@ 4.6 Ablation Study
Ablation Study on Domain Tag Generator: To validate that the proposed domain tag effectively mitigates domain discrepancies and enhances the model’s generalization across diverse data domains, we conducted an ablation study. Specifically, we removed the domain tag generator (DTG) and trained the FakeShield with identical configurations on the MMTD-Set, the test results are displayed in Table 6. Without the DTG, the model’s detection performance declined across test sets from each data domain. Notably, the detection ACC and F1 score for IMD2020 decreased by 0.12 and 0.11. This demonstrates that without the support of the DTG module, the model struggles to effectively differentiate between various data domains, leading to more pronounced data conflicts and a significant reduction in both generalization and practical applicability.
@@
Method CASIA1+ IMD2020 DeepFake AIGC-Editing
ACC F1 ACC F1 ACC F1 ACC F1
Ours w/o DTG 0.92 0.92 0.71 0.79 0.89 0.90 0.72 0.78
Ours 0.95 0.95 0.83 0.90 0.98 0.99 0.93 0.93
@@ A.2 Data Sources
A.3 More Ablation Experiments
Exploring the Introduction of an Error-Correction Mechanism in MFLM: We observe that our MFLM inherently possesses some error-correction capability. This is because, during training, we use potentially inaccurate $\mathbf{O}_{det}$ as input and encourage the model to output $\hat{\mathbf{M}}_{loc}$. Thus, during testing, if $\mathbf{O}_{det}$ contains minor inaccuracies, such as incorrect descriptions of tampered locations, MFLM can correct them. To further explore the impact of incorporating an error correction mechanism during MFLM training on localization accuracy, we conducted an ablation study. During MFLM training, we used the ground truth $\hat{\mathbf{O}}_{det}$ to constrain TCM’s output to correct the input $\mathbf{O}_{det}$. The results in the Table 8 indicate that correcting $\mathbf{O}_{det}$ does not improve $\mathbf{M}_{loc}$ predictions, possibly due to interference between mask and text optimization.
Table 8: Exploration on introducing error correction in MFLM on localization performance.
Method CASIA1+ IMD2020 DeepFake AIGC-Editing
IoU F1 IoU F1 IoU F1 IoU F1
Using correct $\mathbf{O}_{det}$ 0.51 0.56 0.49 0.55 0.13 0.21 0.12 0.15
FakeShield 0.54 0.60 0.50 0.57 0.14 0.22 0.18 0.24
More Ablation Study on LLM in the DTE-FDM: To investigate the roles of $\mathbf{T}_{tag}$, $\mathbf{T}_{img}$, and $\mathbf{T}_{ins}$. When training MFLM, we kept other training configurations unchanged and adjusted the inputs to: $\mathbf{T}_{ins},\mathbf{T}_{img}$} and $\mathbf{T}_{ins},\mathbf{T}_{tag}$}. The test results are shown in Table 9. It is evident that using input $\mathbf{T}_{ins},\mathbf{T}_{img}$} is slightly better than $\mathbf{T}_{ins},\mathbf{T}_{tag}$}, but both are inferior to our method with the input $\mathbf{O}_{det},\mathbf{T}_{img}$}.This indicates that using an LLM to first analyze the basis for tampering and then guiding the localization achieves better results.
Table 9: Ablation Study on LLM in the DTE-FDM with different input combinations.
Method CASIA1+ IMD2020 DeepFake AIGC-Editing
IoU F1 IoU F1 IoU F1 IoU F1
$\mathbf{T}_{ins},\mathbf{T}_{img}$ 0.50 0.55 0.48 0.53 0.13 0.21 0.12 0.15
$\mathbf{T}_{ins},\mathbf{T}_{tag}$ 0.49 0.54 0.47 0.52 0.12 0.19 0.12 0.14
$\mathbf{T}_{ins},\mathbf{T}_{tag},\mathbf{T}_{img}$ 0.51 0.55 0.48 0.54 0.13 0.2 0.11 0.14
$\mathbf{O}_{det},\mathbf{T}_{img}$ (Ours) 0.54 0.60 0.50 0.57 0.14 0.22 0.18 0.24
A.4 More AIGC-Editing generalization experiments
To further verify the model’s generalization performance in the AIGC-Editing data domain, we expanded our test set by constructing 2,000 tampered images using controlnet inpainting (Zhang et al., 2023) and SDXL inpainting (Podell et al., 2023), which the model had not encountered before. We selected MVSS-Net, CAT-Net, and HiFi-Net, which performed well in the AIGC-Editing data domain in Tables 1, as comparison methods for testing. The detection and localization test results are presented in Table 10. Our method demonstrates leading performance on unseen datasets and exhibits strong generalization capabilities.
Table 10: Generalized performance comparison on ControlNet and SDXL Inpainting datasets.
Method ControlNet Inpainting SDXL Inpainting
ACC Image-level F1 IoU Pixel-level F1 ACC Image-level F1 IoU Pixel-level F1
MVSS-Net 0.38 0.27 0.05 0.09 0.35 0.20 0.05 0.08
CAT-Net 0.90 0.89 0.11 0.17 0.86 0.83 0.06 0.09
HiFi-Net 0.49 0.66 0.17 0.28 0.44 0.61 0.02 0.03
Ours 0.99 0.99 0.18 0.24 0.99 0.99 0.18 0.23
A.5 Answer Analysis